2x+3y-z=-1
x-5y-3z=0
3x+4y-1=0

Answer by TimothyLamb(4379)

2x+3y-z=-1
x-5y-3z=0
3x+4y-1=0
---
put the system of linear equations into standard form
---
2x+3y-z=-1
x-5y-3z=0
3x+4y=1
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 1.1818182
y = -0.63636364
z = 1.4545455
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php